OpenStudy (patrickweewoos):

5 medals for five questions!!!!

OpenStudy (steve816):

Okee patrick what is your question

OpenStudy (patrickweewoos):

Give me a second to get them

OpenStudy (dm91):

what are the questions?

OpenStudy (patrickweewoos):

solve 4y=144 y=1

OpenStudy (patrickweewoos):

@steve816

OpenStudy (steve816):

4y=144 144/4=36 That is your answer patrick 36. Now go play with spongebob.

OpenStudy (patrickweewoos):

-1/8c=2

OpenStudy (patrickweewoos):

@steve816

OpenStudy (patrickweewoos):

@steve816 Did you want 5 medals?

OpenStudy (steve816):

-1/16

OpenStudy (steve816):

Stop asking questions patrick.

OpenStudy (patrickweewoos):

-16? @steve816

OpenStudy (patrickweewoos):

The answer choices are: 16 4 -4 -16

OpenStudy (patrickweewoos):

-30=6z @steve816 And which one dude

OpenStudy (steve816):

-5 Now STOP ASKING ME QUESTIONS DUDE!!!

OpenStudy (patrickweewoos):

@steve816 for which one?

OpenStudy (patrickweewoos):

I needed the top one! @steve816 and why be on open study if you dont want questions

OpenStudy (steve816):

JUST STAHP

OpenStudy (patrickweewoos):

well?

OpenStudy (dm91):

patrick you van ask me if you want but i am not that good but i will try

OpenStudy (dm91):

can*

OpenStudy (patrickweewoos):

Well @steve816 would have gotten 5 medals for those 2 more

OpenStudy (dm91):

ya

OpenStudy (patrickweewoos):

here

OpenStudy (dm91):

ok i will try but i am not saying i am good at it

OpenStudy (patrickweewoos):

-5/6n=40

OpenStudy (patrickweewoos):

@steve816

OpenStudy (dm91):

what are the choices?

OpenStudy (patrickweewoos):

56 -56 48 -48

OpenStudy (dm91):

try 48